# Knowledge Adaptation as Posterior Correction

### v1

@@ Abstract
Adaptation is the holy grail of intelligence, but even the best AI models (like GPT) lack the adaptivity of toddlers. So the question remains: how can machines adapt quickly? Despite a lot of progress on model adaptation to facilitate continual and federated learning, as well as model merging, editing, unlearning, etc., little is known about the mechanisms by which machines can naturally learn to adapt in a similar way as humans and animals. Here, we show that all such adaptation methods can be seen as different ways of ‘correcting’ the approximate posteriors. More accurate posteriors lead to smaller corrections, which in turn imply quicker adaptation. The result is obtained by using a dual-perspective of the Bayesian Learning Rule of Khan and Rue (2023) where interference created during adaptation is characterized by the natural-gradient mismatch over the past data. We present many examples to demonstrate the use of posterior-correction as a natural mechanism for the machines to learn to adapt quickly.
@@ 1 Introduction
Quick adaptation is essential for survival in nature so much so that it can be equated to intelligence (Sternberg, 2019). Despite this, there is no Artificial Intelligence (AI) system to this day whose adaptive abilities are considered anywhere close to those of humans and animals. Large Language Models (LLMs) like GPT and Gemini can do many tasks better than humans, but they lack the adaptivity of toddlers. We cannot simply teach such models new things by just talking or interacting with them. Hence, there are no simple ways to quickly fix their harmful behavior reported nowadays on a daily basis. There is no magical tool to just permanently add, remove, or modify the knowledge of LLMs. The most trusted way is to simply retrain the whole system which is too costly to fix their day to day mistakes. So the question remains: how can such models adapt quickly? Quick adaptation is important not only to reduce the cost but also to improve the safety, security, and sustainability of AI.
@@
A lot has been done recently to instill adaptivity in large AI models but efforts are fragmented with no apparent similarities in the methodologies used (Fig. 1). For example, Continual Learning methods adapt to streaming data over time (Kirkpatrick et al., 2017), while Federated Learning methods assimilate knowledge spread across space (McMahan et al., 2017). While both deal with distributed information, fundamentally different methodologies are preferred, for instance, replay of old data is common in continual learning (Rebuffi et al., 2017; Pan et al., 2020; Buzzega et al., 2020) but discouraged in federated learning because of privacy. Such differences can make it harder to connect the two sub-fields at a fundamental level. The same is true of other sub-fields, such as, model merging and editing (Wortsman et al., 2022a, b; Ilharco et al., 2023; Mitchell et al., 2022), knowledge distillation (Hinton et al., 2015), fine-tuning (Hu et al., 2022), etc. All of these efforts use adaptation to avoid retraining as much as possible, but they remain disconnected from each other. Our goal in this this paper is to unify these approaches to unravel a common adaptation mechanism behind them. The underlying mechanism can then be used to not only explain the effectiveness of these approaches but also to find new ways to improve them.
[…]
@@ 2 Posterior Correction
We present a new ‘posterior correction’ approach to unify existing adaptation schemes. The approach relies on a variational reformulation where the goal of learning is to find accurate approximations of the posterior distribution (Khan and Rue, 2023). For example, an Empirical Risk Minimization (ERM) problem over model parameters $\boldsymbol{\theta}\in\Theta$ is reformulated as Variational Learning (VL) over candidate distribution $q(\boldsymbol{\theta})\in\mathcal{Q}$, that is, we reformulate
$\boldsymbol{\theta}_{t}=\arg\min_{\boldsymbol{\theta}\in\Theta}\,\,\sum_{i=0}^{t}\ell_{i}(\boldsymbol{\theta})\qquad\text{as}\qquad q_{t}=\arg\min_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}].$ (1)
@@
Here, ERM uses the loss $\ell_{i}$ defined over $t$ data examples for $i=1,2,\ldots,t$, as well as a regularizer $\ell_{0}$ which can either be added explicitly or implemented implicitly through an algorithm. In contrast, the VL problem uses expectation of losses and a prior $p_{0}\propto\exp(-\ell_{0})$ used in the Kullback-Leibler (KL) divergence term. Despite these differences, the ERM solution $\boldsymbol{\theta}_{t}$ can be recovered from VL by using a Gaussian $q_{t}$ and applying the delta method, as described in Sec. 3. We will use the above reformulation to derive adaptation strategies proposed for ERM as a special case of the posterior-correction approach for VL.
We start with a simple case to set the idea. Suppose we receive a new example $t+1$ with loss $\ell_{t+1}$, then how can we quickly adapt $q_{t}$ to recover $q_{t+1}$? Bayes’ rule suggest to use $q_{t}$ as the prior and update the new posterior as follows: $p_{t+1}\propto e^{-\ell_{t+1}}q_{t}$. However, this is intractable unless the loss and prior form a conjugate pair, like Gaussians (Bishop, 2006, pp. 117). A tractable alternative is to use VL to keep the update within the set $\mathcal{Q}$,
$\hat{q}_{t+1}=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}].$ (2)
@@
Unfortunately, this does not exactly recover the true $q_{t+1}$. We will now show that $q_{t+1}$ can be recovered exactly if we add a correction term in the above update.
[…]
@@
$q_{t}(\boldsymbol{\theta})\propto\prod_{i=0}^{t}\exp\left(-\hat{\ell}_{i|t}(\boldsymbol{\theta})\right),\,\,\text{ where }\hat{\ell}_{i|t}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle.$ (3)
Here, $\widetilde{\nabla}$ denotes the natural gradient with respect to $\boldsymbol{\lambda}$ evaluated at the natural parameter of $q_{t}$. A derivation is given in Sec. 4.1 where we also allow non-constant base measure. [^1] For non-constant base measure $h(\boldsymbol{\theta})$, we just need to modify ^$\hat{\ell}_{0|t}=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{0}+\log h]\rangle-\log h$. Table 1 shows a few examples of the sites ^$\hat{\ell}_{i|t}$ where see Gaussian sites resembles those obtained using Taylor’s method but they use derivatives evaluated and averaged over the samples from $q_{t}$; also see Fig. 2. The site functions contain more global information than Taylor’s surrogates and they also apply more generally, for instance, to discontinuous loss functions and discrete variables. More details are included in Sec. 4.1.
[…]
Form of $q_{t}$ Expression for the site ^$\hat{\ell}_{i|t}$
@@
Figure 2: The left panel compares the site ^$\hat{\ell}^{\text{iso}}_{i|t}$ to the $1^{\text{st}}$-order Taylor of $\ell_{i}$ at $\mathbf{m}_{t}$. The right panel compares ^$\hat{\ell}^{\text{full}}_{i|t}$ to the $2^{\text{nd}}$-order Taylor. The site functions use expectations of the gradients and Hessians over $q_{t}$ and capture a more global information around $\mathbf{m}_{t}$.
We use Eq. 3 to correct the update in Eq. 2. In the first line below, we begin with the definition of $q_{t+1}$, then expand the KL term where we divide and multiply by $q_{t}$ and its dual form respectively, and rearrange terms to get the correction shown in the last line:
$\begin{split}q_{t+1}&=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{E}_{q}\left[\log\left(\frac{q}{e^{-\ell_{0}}}\times\frac{\prod_{i=0}^{t}e^{-\hat{\ell}_{i|t}}}{q_{t}}\right)\right]\\ &=\arg\min_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\sum_{i=0}^{t}\underbrace{\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]}_{\text{Correction}}\end{split}$ (4)
Adding the third term corrects the update in Eq. 2 to ensure the recovery of the exact $q_{t+1}$ (instead of $\hat{q}_{t+1}$). Corrections are required in all the past $\ell_{i}$ as well as the regularizer $\ell_{0}$. Essentially, past $\ell_{i}$ are represented in $q_{t}$ through sites ^$\hat{\ell}_{i|t}$ and, when a new $\ell_{t+1}$ is added, the representation needs to change to ^$\hat{\ell}_{i|t+1}$. The correction term gives a precise mathematical characterization of the interference created during adaptation, which is an age-old problem (Sutton, 1986) without a precise agreed-upon mathematical definition. Other views of posterior corrections are discussed in Sec. 4.2 where it is used to derive a Bayes’ filter and a new definition of information gain, and also interpreted as natural-gradient mismatch. In addition, a method to ‘boost’ training trajectories is also discussed.
The main result of this paper is to show that many existing adaptation methods tailored to handle specific adaptation cases, can all be seen as specific instances of posterior correction. Smaller correction imply less interference which can be handled quickly during adaptation. This key message is demonstrated by covering multiple adaptation scenarios, such as, continual learning, influence estimation, model merging, and federated learning.
@@ 3 Knowledge Adaptation as Posterior Correction
To derive existing methods as posterior correction, we will reformulate an ERM problem as a VL problem. For instance, for a $q^{\text{iso}}$ family, VL in Eq. 1 reduces to an ERM over the mean $\mathbf{m}$ by using the $1^{\text{st}}$-order delta method (Khan and Rue, 2023, App. C),
$\mathbb{E}_{q}[\ell_{i}]\approx\ell_{i}(\mathbf{m})\quad\implies\quad\mathbf{m}_{t}=\arg\min_{\mathbf{m}\in\Theta}\,\,\sum_{i=0}^{t}\ell_{i}(\mathbf{m})+\text{const.}$. (5)
This is because the entropy of $q^{\text{iso}}$ is constant, therefore the KL term in VL only contributes the $\ell_{0}$ term. This reformulation is suitable for ERMs that aim for $1^{\text{st}}$-order stationarity conditions, for instance, by using stochastic gradient descent. For algorithms that aim $2^{\text{nd}}$-order optimality, we can use a $2^{\text{nd}}$-order delta method over a $q^{\text{full}}$ family and assume that the covariance is set to the Hessian at the ERM solution $\mathbf{m}_{t}$ which corresponds to Laplace’s method (Khan and Rue, 2023, App. C.1). Similarly, for methods such as Adam, SAM, or RMSprop, a $q^{\text{diag}}$ family is suitable. The explicit choice to not optimize for the covariance is made in non-Bayesian methods, and by making the same choice in the VL problem, we can recover them by solving a variational formulation.

### v3

@@ Abstract
Adaptation is the holy grail of intelligence, but even the best AI models lack the adaptability of toddlers. In spite of great progress, little is known about the mechanisms by which machines can learn to adapt as fast as humans and animals. Here, we cast adaptation as ‘correction’ of old posteriors and show that a wide-variety of existing adaptation methods follow this very principle, including those used for continual learning, federated learning, unlearning, and model merging. In all these settings, more accurate posteriors often lead to smaller corrections and can enable faster adaptation. Posterior correction is derived by using the dual representation of the Bayesian Learning Rule of khanRue23, where the interference between the old representation and new information is quantified by using the natural-gradient mismatch. We present many examples demonstrating how machines can learn to adapt quickly by using posterior correction.
Keywords: Variational Bayes, Continual Learning, Federated Learning, Model Merging.
@@ 1 Introduction
Quick adaptation is essential for survival in nature so much so that it can be equated to intelligence (sternberg2019theory). To this day, there is no Artificial Intelligence (AI) system whose adaptive abilities are considered anywhere close to those of humans and animals. Large Language Models (LLMs) like GPT and Gemini seem to do many tasks better than experts, but lack adaptability of toddlers. The only sense in which they adapt, through prompting, is temporary and arguably superficial. We cannot teach new things to such models by interacting with them. There are no easy ways to quickly fix their mistakes and no magical tool to permanently add, remove, or modify their knowledge. The most trusted way is to simply retrain the whole system which is too costly to fix problematic behaviors reported on a daily basis.
@@
A lot has been done recently to instill adaptivity in large AI models but these efforts are fragmented with no apparent similarities among the prominent methodologies (Fig. 1). For example, continual-learning methods adapt to streaming data over time (kirkpatrick2017overcoming), while federated-learning methods assimilate knowledge spread across different locations (mcmahan2016fedavg). While both deal with distributed information, they use fundamentally different methodologies. As an example, replay of old data is common in continual learning (rebuffi2017icarl; pan2020continual; buzzega2020dark) but discouraged in federated learning due to privacy concerns. Such differences can make it harder to connect these two fields at a fundamental level. The same is true of other approaches, such as, model merging (wortsman22modelsoups; wortsman2021robust; IlharcoRWSHF23; DBLP:conf/iclr/MitchellLBFM22), knowledge distillation (hinton2015distilling), fine-tuning (DBLP:conf/iclr/HuSWALWWC22), etc. All such approaches attempt to adapt and reuse old knowledge quickly and avoid retraining as much as possible, but they remain disconnected from each other.
Our goal in this paper is to unify these efforts and unravel a common adaptation mechanism behind them. We propose to formulate adaptation of old knowledge as ‘correction’ of old posterior approximations. For example, in continual learning, inclusion of new data can be expressed as matching the predictions of the new and old posteriors. We give several such examples where knowledge adaptation, when formulated as posterior correction, amounts to matching the predictions, gradients, and/or Hessians. Through examples in problems such as continual learning, unlearning, model merging, and federated learning, we argue that better posterior approximations reduce the correction needed to adapt and can facilitate faster adaptation. The proposal can also be useful for traditional approaches such as domain adaptation (ben2006analysis) and semi-supervised learning (olivier2006semi), although we focus on recent efforts to instill adaptivity in large AI models. Ultimately, posterior correction provides a new way to form and inject prior knowledge for faster adaptation, even in cases where there are no Bayesian models available or exact computation of the posterior is infeasible.
Figure 1: Four popular scenarios for adaptation of model parameters $\boldsymbol{\theta}$. Black arrows indicate the flow of adapted knowledge, while the gray arrows indicate pre-trained knowledge. Continual learning adapts $\boldsymbol{\theta}_{t}$ to $\boldsymbol{\theta}_{t+1}$ to include new data $\mbox{${\cal D}$}_{t+1}$. Unlearning aims to estimate the model after removing a specific data set $\mbox{${\cal D}$}_{i}$. Model merging attempts to improve a pre-trained base model $\boldsymbol{\theta}_{0}$ by merging back the fine-tuned models $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$. Finally, federated learning aims to obtain a joint model $\boldsymbol{\theta}_{\text{jnt}}$ by using locally trained models.
@@ 2 Posterior Correction
To derive posterior correction, we will use a variational-Bayesian framework that applies to a wide variety of learning problems, including Bayesian and non-Bayesian approaches. The goal of variational learning (VL) is to find good approximations of the posterior distribution (khanRue23), but it can also be seen as a generalization of the well-known Empirical Risk Minimization (ERM) problem. For instance, the ERM problem of optimizing over model parameters $\boldsymbol{\theta}\in\Theta$ can also be reformulated as a VL problem where we optimize over the distributions $q(\boldsymbol{\theta})\in\mathcal{Q}$. That is, we can write
$\boldsymbol{\theta}_{t}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\Theta}\,\,\sum_{i=0}^{t}\ell_{i}(\boldsymbol{\theta})\qquad\text{as}\qquad q_{t}=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}].$ (1)
@@
The ERM objective on the left uses loss functions $\ell_{i}$ defined over data sets $\mbox{${\cal D}$}_{i}$ indexed by $i=1,2,\ldots,t$, and we denote the regularizer by $\ell_{0}$ and it can either be added explicitly, or obtained implicitly through a learning algorithm. In contrast to ERM, the VL problem uses the expected loss $\mathbb{E}_{q}[\ell_{i}]$ under $q(\boldsymbol{\theta})$ and a prior $p_{0}\propto\exp(-\ell_{0})$ in the Kullback-Leibler (KL) divergence term. Despite these differences, as shown in Sec. 3, the ERM solution $\boldsymbol{\theta}_{t}$ can be recovered as a special case of a VL problem in which $\mathcal{Q}$ is the set of Gaussian distributions. In this case, the mean $\mathbf{m}_{t}$ of the Gaussian $q_{t}$ is equal to $\boldsymbol{\theta}_{t}$ when we use the delta method to approximate the expectations $\mathbb{E}_{q_{t}}[\ell_{i}]\approx\ell_{i}(\mathbf{m}_{t})$ in the VL objective.
Posterior correction uses the VL framework to cast knowledge adaptation as correction of old posteriors. We will now demonstrate the main idea for the problem of continual learning where the goal is to adapt to a new task $t+1$ an old model trained on $t$ previous tasks. In VL terms, this goal translates to the adaptation of the posterior $q_{t}$ shown in Eq. 1 to obtain the posterior $q_{t+1}$, which additionally includes the loss on the new task (denoted by $\ell_{t+1}$). The optimal update is to use the Bayes’ rule, that is, we compute the new posterior $p_{t+1}$ as follows: $p_{t+1}\propto q_{t}\exp\left(-\ell_{t+1}\right)$ where $q_{t}$ is the prior and $\ell_{t+1}$ defines the likelihood. However, this is challenging unless the loss and prior form a conjugate pair, like Gaussians (bishop2006pattern, p. 117). A tractable alternative is Variational Continual Learning (cuong2018), which uses the following recursive variational form setting the prior to $q_{t}$,
$\hat{q}_{t+1}=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}].$ (2)
@@
Unfortunately, this does not exactly recover the true $q_{t+1}$. Soon we will see that exact recovery is possible if we include a correction term in the above update.
Figure 2: We illustrate the dual form given in Eq. 3 on a 1-D binary classification example (circle vs square in the leftmost panel). The classifier is simply a threshold $\theta$, fit via a logistic likelihood (pink sigmoid). We use a Gaussian $q_{t}(\theta)$ with mean $m_{t}$ shown with vertical dashed line. The logistic losses for the three examples are shown in the top row of middle panel (transparent lines) along with their quadratic sites (solid lines). The bottom row displays likelihoods. The rightmost panel applies Eq. 3 to form $q_{t}$. The example farthest from the classifier has the smallest curvature and least contribution to $q_{t}$. The example is discussed in detail in Sec. A.1 along with a general case of logistic regression in Sec. A.2.
The correction term can be derived by using a dual representation of $q_{t}$, originally proposed by khan2018fast1. Essentially, we can write $q_{t}$ as the product of local factors defined using site functions for each $\ell_{i}$. The sites are essentially surrogates of the loss which often take a simple form; see an illustration in Fig. 2. More precisely, let $\mathcal{Q}$ be the set of minimal exponential-family distributions $q(\boldsymbol{\theta})=\exp(\langle\mathbf{T}(\boldsymbol{\theta}),\boldsymbol{\lambda}\rangle)/\mathcal{Z}$ with sufficient statistics $\mathbf{T}(\boldsymbol{\theta})$, natural parameter $\boldsymbol{\lambda}$, and partition function $\mathcal{Z}$; the inner product is denoted by $\langle\cdot,\cdot\rangle$. Then, a solution $q_{t}$ of Eq. 1 can be written as a product of local factors as follows,
@@
$q_{t}(\boldsymbol{\theta})=\frac{1}{\mathcal{Z}_{t}}\prod_{i=0}^{t}\exp\left(-\hat{\ell}_{i|t}(\boldsymbol{\theta})\right),\,\,\text{ where }\hat{\ell}_{i|t}(\boldsymbol{\theta})=\langle\mathbf{T}(\boldsymbol{\theta}),\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]\rangle.$ (3)
By $\hat{\ell}_{i|t}$ we denote the site function for loss $\ell_{i}$ evaluated at $q_{t}$. The expression of the site function shows that they are obtained by using natural gradients, denoted by $\smash{\widetilde{\nabla}\mathbb{E}_{q_{t}}[\ell_{i}]}$, taken with respect to the natural parameter $\boldsymbol{\lambda}_{t}$ of $q_{t}$. A definition of natural gradients and detailed derivation of the above expression are both provided in Sec. 4.1, along with a generalization to the iterates of the Bayesian learning rule (BLR) (khanRue23) which is later used to boost learning trajectories.
Table 1 lists some examples of the sites which will be used later in the paper. We note that, for Gaussian posteriors, the sites resemble Taylor’s approximations of the loss. The main difference is that the sites use derivatives evaluated and averaged over samples from $q_{t}$; see Fig. 3. The site functions contain more global information than Taylor surrogates and they apply more generally, for instance, to discontinuous losses over discrete variables.
Table 1: Site functions. We denote $\bar{\mathbf{H}}_{i|t}=\mathbb{E}_{q_{t}}[\nabla^{2}\ell_{i}]$ and its diagonal by the vector $\bar{}\mbox{$\mbox{$\mathbf{h}$}$}_{i}$. Element-wise multiplication and divisions are denoted by $\mbox{$\mbox{$\mathbf{a}$}$}\cdot\mbox{$\mbox{$\mathbf{b}$}$}$ and $\mbox{$\mbox{$\mathbf{a}$}$}/\mbox{$\mbox{$\mathbf{b}$}$}$ respectively.
Form of $q_{t}$ Expression for the site $\hat{\ell}_{i|t}$
@@
Figure 3: The left panel compares the site $\hat{\ell}^{\text{iso}}_{i|t}$ to the first-order Taylor expansion of $\ell_{i}$ at $\mathbf{m}_{t}$. The right panel compares $\hat{\ell}^{\text{full}}_{i|t}$ to the second-order Taylor. The sites use expectations of the gradients and Hessians over $q_{t}$ and capture more global information around $\mathbf{m}_{t}$.
We use Eq. 3 to correct the update in Eq. 2 to recover $q_{t+1}$ exactly, as shown in the equation below. In the first line, we begin with the definition of $q_{t+1}$ from Eq. 1. In the second line, we expand the KL term and divide and multiply by $q_{t}$. In the third line, we rewrite the $q_{t}$ that appears in the numerator in its dual form, and finally we rearrange terms to get the correction term in the last line,
$\begin{split}q_{t+1}&=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\sum_{i=1}^{t+1}\mathbb{E}_{q}[\ell_{i}]+\mathbb{D}_{\text{KL}}[q\,\|\,p_{0}]\\ &=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{E}_{q}\left[\log\left(\frac{q}{p_{0}}\times\frac{q_{t}}{q_{t}}\right)\right]\\ &=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\sum_{i=1}^{t}\mathbb{E}_{q}[\ell_{i}]+\mathbb{E}_{q}\left[\log\left(\frac{q}{\frac{1}{\mathcal{Z}_{0}}e^{-\ell_{0}}}\times\frac{\frac{1}{\mathcal{Z}_{t}}\prod_{i=0}^{t}e^{-\hat{\ell}_{i|t}}}{q_{t}}\right)\right]\\ &=\operatorname*{arg\,min}_{q\in\mathcal{Q}}\,\,\mathbb{E}_{q}[\ell_{t+1}]+\mathbb{D}_{\text{KL}}[q\,\|\,q_{t}]+\sum_{i=0}^{t}\underbrace{\mathbb{E}_{q}[\ell_{i}-\hat{\ell}_{i|t}]}_{\text{Correction}}\end{split}$ (4)
The third term in the last line corrects the update of Eq. 2, ensuring the exact recovery of $q_{t+1}$ (instead of the approximate $\hat{q}_{t+1}$). Corrections are required for all the past $\ell_{i}$’s as well as for the regularizer $\ell_{0}$. Essentially, past $\ell_{i}$’s are represented in $q_{t}$ through sites $\hat{\ell}_{i|t}$ and, when a new $\ell_{t+1}$ is added, the site $\hat{\ell}_{i|t}$ needs to change to $\hat{\ell}_{i|t+1}$. The correction term gives a precise mathematical characterization of this interference between the old representation and new information, which is an age-old problem (sutton1986two) without a precise agreed-upon mathematical definition. Other views of posterior correction are discussed in Sec. 4.2 where the correction is expressed as a Bayesian filter, a new definition of information gain is given, and interference is interpreted as the natural-gradient mismatch computed over the past data. In addition, we discuss a method to boost training trajectories.
@@ 3 Knowledge Adaptation as Posterior Correction
The main result of this paper is to show that many existing adaptation methods tailored to handle specific adaptation cases, can all be seen as specific instances of posterior correction. This key message is demonstrated by covering multiple adaptation scenarios, such as, continual learning, unlearning, model merging, and federated learning. The results are derived by relying on the following three types of Gaussian families defined in Table 1:
  1. 1.
    The ‘isotropic Gaussian’ family is denoted by candidates $q^{\text{iso}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m},\mbox{$\mbox{$\mathbf{I}$}$})$ where the mean $\mathbf{m}$ is learned but the covariance is fixed to the identity matrix $\mathbf{I}$;
  2. 2.
    The ‘diagonal Gaussian’ family denoted by candidates $q^{\text{diag}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m},\text{diag}(\mbox{$\mbox{$\mathbf{s}$}$})^{-1})$ where the mean $\mathbf{m}$ and the diagonal vector $\mathbf{s}$ of the precision matrix are both learned;
  3. 3.
    The ‘full Gaussian’ family denoted by $q^{\text{full}}=\mbox{${\cal N}$}(\boldsymbol{\theta}|\mathbf{m},\mbox{$\mbox{$\mathbf{S}$}$}^{-1})$ where $\mathbf{m}$ and $\mathbf{S}$ are learned.
We will see that increasing the size of the family from $q^{\text{iso}}$ to $q^{\text{diag}}$ and then to $q^{\text{full}}$ often leads to reductions in corrections. Smaller corrections imply less interference and yield better adaptation methods. This message is repeated throughout the paper through many examples, all aimed at establishing posterior-correction as a unifying principle for knowledge adaptation. We do not derive a precise relation between the size of corrections and speed of adaptation. That would require in-depth studies for each case separately which is out of the scope for this work. Our hope is that the future work will benefit from the insights of this paper to perform such studies for different types of adaptation tasks.
Throughout, we will reformulate an ERM problem as a VL problem. For instance, for a $q^{\text{iso}}$ family, VL in Eq. 1 reduces to an ERM over the mean $\mathbf{m}$ by using the first-order delta method (khanRue23, App. C),
$\mathbb{E}_{q}[\ell_{i}]\approx\ell_{i}(\mathbf{m})\quad\implies\quad\mathbf{m}_{t}=\operatorname*{arg\,min}_{\mathbf{m}\in\Theta}\,\,\sum_{i=0}^{t}\ell_{i}(\mathbf{m})+\text{const.}$ (5)
This is because the entropy of $q^{\text{iso}}$ is constant, therefore the KL term in VL only contributes the $\ell_{0}$ term. This reformulation is suitable for ERMs that aim for the first-order stationarity condition, for instance, by using stochastic gradient descent. For algorithms that aim for second-order optimality, we can use a second-order delta method over a $q^{\text{full}}$ family and assume that the covariance is set to the Hessian at the ERM solution $\mathbf{m}_{t}$ which corresponds to Laplace’s method (khanRue23, App. C.1). We start with the first adaptation task of continual learning in the next section.